Grade 11Magical Mathematics[Interesting Approach]

A train started from station 'A' and proceeded towards station 'B' at a speed of 48 Km/hr. Forty-five minutes later another train started from station 'B' and proceeded towards station 'A' at 50 Km/hr. If the distance between the two stations is 232 Km, at what distance from station 'A' will the trains meet?

a)

doppalapudi sairam
15 Years agoGrade 11

1 Answer

AKASH GOYAL AskiitiansExpert-IITD
15 Years ago

Dear Student

They will meet when sum of the distance covered by them is 232 km

suppose they meet after t hr of departure time at A

then travelling time of B will be t-3/4 hrs

DA=48t

DB=50(t-3/4)

DA +DB=  48t + 50(t-3/4)=232

solving we get t=11/4 hrs

distance from A= 48x11/4=132 km

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..